# Exploiting Sparsity for Long Context Inference: Million Token Contexts on Commodity GPUs

### v1

@@ Abstract
There is growing demand for performing inference with hundreds of thousands of input tokens on trained transformer models. Inference at this extreme scale demands significant computational resources, hindering the application of transformers at long contexts on commodity (i.e not data center scale) hardware. To address the inference time costs associated with running self-attention based transformer language models on long contexts and enable their adoption on widely available hardware, we propose a tunable mechanism that reduces the cost of the forward pass by attending to only the most relevant tokens at every generation step using a top-k selection mechanism. We showcase the efficiency gains afforded by our method by performing inference on context windows up to 1M tokens using approximately 16GB of GPU RAM. Our experiments reveal that models are capable of handling the sparsity induced by the reduced number of keys and values. By attending to less than 2% of input tokens, we achieve over 95% of model performance on common long context benchmarks (LM-Eval, AlpacaEval, and RULER).
Machine Learning, ICML
@@ 1 Introduction
Figure 1: (top) Typical attention requires each query vector to compute an inner product with each key vector in the context window. In practice, most key vectors are insignificant to attention, so much of this computation is wasted. (left-bottom) top-k attention retrieves only the keys that contribute significantly to the attention computation, leaving the gray arrows out and achieving sublinear runtime.
@@
The prefill stage incurs an $O(N^{2})$ compute and memory cost due to the large attention matrix formed in the computation (where $N$ is the number of tokens in the context). To mitigate this, brute-force methods such as Ring Attention (Liu et al., 2023) have been developed. This uses round-robin communication between servers to scale computation linearly in the number of GPUs. Importantly, the prefill stage occurs only once, while the decoding stage occurs many hundreds of times.
@@
Cache decoding incurs an $O(N)$ compute and memory cost, as each new token must attend over all previous tokens in the context. As context lengths grow, the KV cache can become prohibitively large. To see this, note that each token in the context is embedded into a vector of size $D$, where $D$ is the hidden dimension. The number of floating point numbers in the cache is then $2NDL$, where $L$ is the number of network layers. For a value of $D=4096$ and $L=32$ as in (Grattafiori et al., 2024), at $N=100,000$ the memory required for the cache alone is 52GB, assuming 16 bit floating point format (Zhang et al., 2024). At this length the cache is unable to fit on most commodity GPUs.
@@
Offloading the KV cache to the CPU can save memory, but popular offloading strategies such as (Sheng et al., 2023) move all vectors in the cache for each layer at a time. This can cause severe data movement costs, as one layer’s worth of data is 1.6GB as in our above example. These costs accumulate as decoding occurs hundreds or even thousands of times. Cache eviction methods seek to maintain a fixed cache size on the GPU by selectively removing token vectors deemed irrelevant (Zhang et al., 2023), but these strategies ultimately hurt performance as it can be difficult to tell which vectors will be needed in the future before evicting them.
We observe that modern LLMs only require the top contributing tokens to each attention computation to perform well, in line with previous work (Gupta et al., 2021). We exploit this fact to perform fast long context decoding with very little GPU memory. We build an implementation of attention in which keys and values are stored in a vector database in CPU memory. When attention is computed using a query vector at decoding time, we retrieve only the keys with the highest attention scores. This can be done in sublinear time using an approximate k-nearest neighbor search over the cache database, enabling long context inference using plentiful and cheap CPU memory, and without the heavy computational overhead required for full attention. We incur minimal data movement costs since only a fraction of the tokens ever need to be transported on the PCIe bus.
The value of k can vary across layers: some layers may need a larger k budget than others to accurately capture their attention distributions. This allows us to achieve higher accuracy for a fixed computational budget. In addition, we analyze the validity of the top-k assumption through various experiments. While methods involving the restriction of the attention mechanism to the top-k attention scores have been studied, (Gupta et al., 2021; Liu et al., 2024a; Zhang et al., 2024), top-k has yet to be framed in the context of resource-constrained long context decoding with approximate nearest neighbor search. Existing implementations have not been taken to the extreme scale of 1 million tokens.
@@
Figure 2: Performance on OpenLLM Leaderboard (Aidar Myrzakhan, 2024) using only the top-k keys for each attention computation. Typical questions have a context length of $\sim 1000,$ yet only 10 keys are needed to achieve the same performance as full attention.
@@
  • We show that this technique achieves high fidelity on common benchmarks and long-context tasks.
  • We provide an empirical analysis on why top-k attention works and demonstrate its effectiveness on the million token scale.
  • We improve efficiency by varying the k budget on a per-layer basis.
2 Motivation
We observe that modern language models naturally have sparse attention patterns in which a very small number of tokens make up the vast majority of attention mass.
@@ 2 Motivation
We visualize the sparsity of attention in Llama-3-8B using 50 Wikipedia articles in Figure 3. For the last token in each 4000-token context window, we tabulate the number of keys in the context that are needed to collect 75% of the attention mass. Only a small number of the 4000 tokens are needed to collect this mass, especially for deeper layers of the networks.
@@
Next, we take our multi-document long context samples and ask Llama to copy one of the documents that relates to a specific topic. Figure 9 demonstrates that in expectation the attention scores (across all heads and layers) are higher for the document containing the correct document.
Figure 4: Where does the attention go across multiple documents? Observing all of the attention scores over all of the heads and all of the layers we see that in expectation the model is able focus most of its attention on the correct document.
@@
To further analyze the sparsity of the model, we observe the entropy of the attention weights for each token as a function of the layer depth. Figure 5 shows that the entropy is low in all layers and decreases significantly after the first layer. Further analysis of attention sparsity across task categories and intuition on the connection between sparsity and entropy is contained in A.1.
@@
Figure 5: Entropy of the distribution of attention scores for each layer of a model, calculated as $E=-\sum_{i=1}^{N}a_{i}\log(a_{i})$, where $(a_{1},\dots,a_{N})$ is the attention score distribution. Attention score distributions are derived from 50 samples and aggregated over all heads for a given layer. Entropy serves as a measure of how concentrated the attention scores are for a given query token: low entropy indicates a large amount of attention centered over few tokens, and high entropy indicates a more uniform dispersion of attention scores.
@@
The sparsity in Llama’s attention weights suggests that only a few key-value pairs should be needed for the model to perform close to its full capabilities. We test this hypothesis in Figure 2 by evaluating models from the Llama family across varying quantities of key-value pairs. In these experiments, we use the same number of keys for each layer of the network and average the model performance over all tasks on the OpenLLM leaderboard. We see that all three models saturate in performance by the 15th key, and the most recent (and most overtrained) variants of the model require only 10 top keys to achieve the same performance as full-scale attentions. While Figure 3 indicates that there is some spreading of the attention mass for layer 1 of the network, this tail mass seems to be unnecessary for good performance on benchmarks.
These experiments are the basis of our idea of exploiting model sparsity to speed up attention computations. Empirically, very few keys are needed to perform inference, but it is important to select keys that are primary contributors to the attention computation. To exploit this sparsity, we use a vector database to retrieve the top-k most influential tokens, enabling the GPU to perform matrix computations while the keys and values live in CPU memory. This approach alleviates both the computational and memory barriers to doing long context inference on a small GPU.
@@ 3 Methodology
We elaborate on our method for reducing memory costs during inference time using top-k attention. We also introduce some notation and terminology related to inference in transformers.
@@ 3.2 Top-k Attention: Accelerating Decoding
The $V_{\ell}[I]$ and vals are then moved to the GPU. Using these, we perform the final attention computation, combining both the weighted values from the context as well as the value vectors from previously generated tokens. This method gives us a peak GPU memory cost of $O(k)$ GPU memory cost, as opposed to $O(N)$. We find that $k$ can be chosen to be a small fraction (around 1% for some tasks) of $N$ while still recovering near-equivalent performance.
The memory savings this method allows for are highlighted in figure 10.
@@ 3.3 Prefilling a KV Cache at the Million Token Scale
In our top-k experiments, we utilized flash attention and an h100 GPU to prefill the caches for our $N$=1M experiments. This allowed us to generate exact prefilled KV caches with only small modifications to the network to accommodate the extreme memory requirements. We employ a chunking strategy similar to (Gupta et al., 2021). The exact modifications we made in order to prefill caches can be found in the appendix. We note that our experiments mimic an important use case of our method, namely a user with a large amount of documents that owns a limited amount of compute. In this case, the user can rent the required compute on the cloud for prefill once, then make as many queries as they want quickly with our method on their own hardware.
@@ 4 Evaluating Top-k Attention at Scale
We evaluate top-k to highlight the relationship between different values of k and performance. We find that models of various sizes and generations perform well even at small k. In general, we observe that a k equal to 1% of the total length of the context is sufficient to achieve 95% of the performance achieved with full, standard attention.
4.1 Evaluations
We analyze the effectiveness of the top-k attention mechanism across three benchmarks: AlpacaEval, Open LLM Leaderboard v1 tasks, and RULER. Each of these benchmark highlight a different quality of an LLM. RULER tests the model’s long context capability. AlpacaEval measures generation quality of the model. Open LLM Leaderboard v1 tasks test the model’s capabilities such as knowledge. Measuring performance across these datasets presents a comprehensive understanding of top-k.
@@ 4.1 Evaluations
We benchmark top-k attention on AlpacaEval (Dubois et al., 2024). AlpacaEval $2.0$ requires a model to generate responses to $805$ queries. These responses are then compared by an LLM-as-a-Judge with GPT-4 Turbo responses generated from the same query set. The winrate percentage is reported, and the LLM-as-a-Judge is GPT-4 Turbo.
Open LLM Leaderboard Tasks
We investigate the performance of top-k on Open LLM Leaderboard tasks. Particularly, we evaluate different models on various values of k on the average of MMLU (Hendrycks et al., 2020), ARC tasks both easy and challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), winogrande (Keisuke et al., 2019), OpenbookQA (Mihaylov et al., 2018), BoolQ (Clark et al., 2019), and PiQA (Bisk et al., 2020). For each task, we record the normalized accuracy when available; otherwise, we record accuracy. We report the average over tasks. We evaluate the following models on these benchmarks: Llama-1 (7B), Llama-2 (7B), Llama-3 (8B), Llama-3.1 (8B), Vicuna-v1.3 (7B), Llama-2 chat (7B), Llama-3 Instruct (8B), Llama-3.1 instruct (8B), Llama-3.2 1B instruct, and Llama-3.2 3B instruct (Touvron et al., 2023a; Zheng et al., 2023; Touvron et al., 2023b; Dubey et al., 2024; AI, 2024). The experiments are conducted using lm-eval-harness in a zero-shot setting (Gao et al., 2024).
4.2 top-k is Effective at Low k
@@
Top-k Performance on RULER
We evaluate RULER using GradientAI’s Llama-3-8B model that has a trained context length of 262k. We find that very small values of k are sufficient to recover near-baseline performance. For every context length evaluated, $95\%$ of the baseline performance can always be achieved with a k value of $1\%$ or less of the total length. In Table 1, at $k=2$, greater than $60\%$ performance is achieved at all context lengths. The performance on RULER improves as k increases. Nevertheless, even at a context length of $131$k tokens to achieve $\sim 98\%$ performance of the full attention only $12.5\%$ of the attention scores are required. This highlights the effectiveness of top-k on long context.
Interestingly, most mechanical tasks have high success rates with little variation across runs. Conversely, the question-answering tasks (from SQuAD (Rajpurkar et al., 2016) and HotpotQA (Yang et al., 2018)) end up being the tasks most indicative of model capability, with a consistently decreasing score as context length increases and as k decreases. These QA tasks were both the most revealing and had the highest variance across test runs, so allocating an extra compute budget to this subset of tasks would be helpful.
@@
Of the three benchmarks evaluated, AlpacaEval required the largest k as a percentage of context length to achieve $95\%$ of baseline performance, with $2.5\%$ of the context length being required. The trend of small k values achieving near-baseline performance was repeated across model sizes and generations.
Top-k Performance on OpenLLM Leaderboard Tasks
We evaluated various models to find that top-k behavior exists regardless of instruction tuning, model size, or the number of tokens the model was trained on. When comparing the models with different amounts of tokens trained, Table 6 left shows all models exhibit a similar curve with performance quickly saturating by a k value $<10$, regardless of the number of tokens on which the model was trained. Furthermore, when comparing Table 6 left and center, we see that instruction models and pretrained base models exhibit the same behavior, very quickly saturating. Additionally, regardless of model size, in Table 6 right, we see that whether a model is $1$B, $3$B, or $8$B the behavior of top-k is the same.
@@
Figure 7: 1 million token needle-in-a-haystack. Comparison between the cache eviction method of (Xiao et al., 2023) and our method on a needle-in-a-haystack task. The red cells show that attention sinks are incapable of retrieving tokens outside of the local window or early sink tokens. Our method is able to do so with $k=10$ and extends to over 1 million tokens
@@
To demonstrate the scaling that top-k attention permits, we use our method to generate tokens from a context window of one million tokens. We choose RULER’s Needle In A Haystack task to use for the context, and we use GradientAI’s Llama-3-8B model that has a trained context length of 1M (Gradient Team, 2024). We run this on a single-GPU node with a Faiss vector database containing the KV cache. [^1] We used a variety of k values for this experiment and it turned out that k =1 is sufficient to solve Needle In A Haystack with a context length of 1 million tokens. We also compare our method against (Xiao et al., 2023) for needle in a haystack in figure 7.
4.4 Layer and Task Dependence on k
We examine the dependence of k on both tasks and also across layers of the network.
Table 3 shows how the value of k necessary to achieve 95% of base model performance varies across the tasks of RULER. For the needle in a haystack task, only a tiny fraction of the entire context is necessary to achieve 95% of the base performance of the model, whereas nearly 9% is necessary in the case of Word Counting. Most of the tasks fall under the 1% line.
Table 2: Percent of attention scores required to reach 95% of dense attention performance for different categories of tasks from the RULER benchmark. k requirements were measured by performance on contexts from 8,000-128,000 tokens. All samples from Llama3-8B.

### v2

@@ Abstract
There is growing demand for performing inference with hundreds of thousands of input tokens on trained transformer models. Inference at this extreme scale demands significant computational resources, hindering the application of transformers at long contexts on commodity (i.e not data center scale) hardware. To address the inference time costs associated with running self-attention based transformer language models on long contexts and enable their adoption on widely available hardware, we propose a tunable mechanism that reduces the cost of the forward pass by attending to only the most relevant tokens at every generation step using a top-k selection mechanism. We showcase the efficiency gains afforded by our method by performing inference on context windows up to 1M tokens using approximately 16GB of GPU RAM. Our experiments reveal that models are capable of handling the sparsity induced by the reduced number of keys and values. By attending to less than 2% of input tokens, we achieve over 95% of model performance on common benchmarks (RULER, AlpacaEval, and Open LLM Leaderboard). The code is available at https://github.com/ryansynk/topk-decoding.
Machine Learning, Long Context
@@ 1 Introduction
Figure 1: (top) Typical attention requires each query vector to compute an inner product with each key vector in the context window. In practice, most key vectors produce insignificant attention scores, and therefore contribute very little to subsequent hidden states, so much of this computation is wasted. (left-bottom) top-k attention retrieves only the keys that contribute significantly to the attention computation, leaving the gray arrows out and achieving sublinear runtime.
@@
The prefill stage incurs an $O(N^{2})$ compute and memory cost due to the large attention matrix formed in the computation (where $N$ is the number of tokens in the context). To mitigate this, brute-force methods such as Ring Attention (Liu et al., 2023) were developed. This approach uses round-robin communication between servers to scale computation linearly in the number of GPUs. Despite its apparent effectiveness, while the prefill stage occurs only once, because the decoding stage occurs many hundreds of times for a single user request, the algorithm limited in its practicality outside of the data-center.
@@
Every step during the decoding stage incurs an $O(N)$ compute and memory cost, as each new token must attend over all previous tokens in the cached context. As context lengths grow, the KV cache can become prohibitively large. If each of the $N$ tokens in the context is embedded into a vector of size $D$ (the hidden dimension) then the number of floating point numbers in the cache is $2NDL$, where $L$ is the number of network layers. For a value of $D=4096$ and $L=32$ as in the Llama-3 8B architecture (Grattafiori et al., 2024), at $N=100,000$ the memory required for the cache alone is 52GB, assuming 16 bit floating point format (Zhang et al., 2024). At this context length the cache is unable to fit on most commodity GPUs.
@@
One approach to save memory during decoding is to offload the KV cache to CPU memory (Sheng et al., 2023). However, the additional data movement required to implement this strategy is itself prohibitive. In the above example, just a single layer’s worth of KV cache data is 1.6GB and this payload must be schlepped back and forth hundreds or thousands of times during a single generation. Attacking the problem from a different angle, cache eviction methods try to maintain a fixed cache size on the GPU by selectively removing token vectors deemed irrelevant (Zhang et al., 2023), but because it can be difficult to tell which vectors will be needed in the future, these strategies ultimately hurt performance.
@@
Figure 2: Performance on selected OpenLLM Leaderboard tasks using only the top-k keys for each attention computation. Typical questions have a context length of $\sim 1000,$ yet only 10 keys are needed to achieve the same performance as full attention. We evaluate an extended set of models in Figure 6.
In contrast to the compute and communication intensive solutions discussed thus far, we base our proposal on a simple fact:
Modern LLMs only need to pay attention
to a handful of tokens at a time.
We exploit this observation to perform fast long context decoding with very little GPU memory. We build an implementation of attention in which keys and values are stored in a vector database in CPU memory. When attention is computed using a query vector at decoding time, we retrieve only the k keys with the highest attention scores. This can be done in sublinear time using an approximate k-nearest neighbor search over the cache database, enabling long context inference using plentiful and cheap CPU memory, and without the heavy computational overhead required for full attention. Due to the fact that the constraint on the number of relevant keys constrains the CPU-GPU transfer volume, our approach incurs minimal data movement costs compared to other KV cache offloading methods since only a fraction of the token representations ever need to be transported between GPU and CPU memory.
To build a deeper understanding of why our approach works, through an array of controlled experiments we interrogate the core “top-k” assumption on which we base our method; that attention in modern LLMs is an inherently sparse operation. While methods exploring the restriction of the attention mechanism to the top-k scores have been studied extensively (Gupta et al., 2021; Liu et al., 2024a; Zhang et al., 2024), prior work does not specifically endeavor to solve resource constrained long-context decoding using scalable tools like approximate nearest neighbor search, nor do they demonstrate million token context inference on single, commodity GPUs.
@@
  • We show that this technique achieves high fidelity on common benchmarks and long-context tasks (Figure 2).
  • We provide an empirical analysis on why top-k attention works and demonstrate its effectiveness at the million token scale.
  • We demonstrate the flexibility of our method by varying the top-k budget on a per-layer basis.
@@
2 Motivation
The core assumption underpinning our proposal is the fact that modern language models naturally exhibit sparse attention patterns in which a very small number of tokens make up the vast majority of attention mass. We begin by verifying that this is in fact true for popular models in simple settings.
Observing attention sparsity in the wild.
@@ 2 Motivation
In Figure 3, we analyze 50 4000-token text samples consisting of concatenated Wikipedia article snippets. We encode these samples using a forward pass through Llama-3-8B, and visualize the resulting attention scores. For the last token in each context window, we tabulate the number of keys in the context that are needed to collect 75% of the attention mass. Only a small number of the 4000 tokens are needed to collect this mass, especially for deeper layers of the networks.
@@
Next, we take these multi-article samples and prompt the model to copy one of the articles that relates to a specific topic. Figure 4 demonstrates that in expectation the attention scores (across all heads and layers) are higher for the tokens contained within the correct document.
Figure 4: Average attention score per token in a given Wikipedia article for a multi-article long context sample (across all heads and layers). The bar for the article prompted to be copied is highlighted in red. Note how the model focuses its attention on the correct document.
@@
Having confirmed that attention scores are indeed concentrated in terms of both how many tokens are relevant, and which tokens are highlighted in an input sequence, we zoom out and probe for structure in the attention scores in aggregate by quantifying their overall entropy. In Figure 5 we visualize the entropy of the last-token attention scores in a sequence as a function of layer depth. We observe that the entropy is low in all layers and decreases significantly after the first layer. Further analysis of attention sparsity across task categories and intuition on the connection between sparsity and entropy is provided in Section A.1.
@@
Figure 5: Entropy of the distribution of attention scores for each layer of a model, calculated as $E=-\sum_{i=1}^{N}a_{i}\log(a_{i})$, where $(a_{1},\dots,a_{N})$ is the attention score distribution. Attention score distributions are derived from last token of 50 1000-token samples and aggregated over all heads for a given layer. Entropy serves as a measure of how concentrated the attention scores are for a given query token: low entropy indicates a large amount of attention centered over few tokens, and high entropy indicates a more uniform dispersion of attention scores. Maximum entropy occurs when the distribution is perfectly uniform, and for 1000-token contexts is $-\log(\frac{1}{1000})$.
Is it safe to exploit attention sparsity?
The observations made in this series of small scale experiments together suggest that only a few key-value pairs should be needed for the model to perform close to its full capabilities. To demonstrate that the observed sparsity can actually be leveraged without compromising performance, we perform a more realistic test.
@@
We evaluate models from the Llama family on small selection of knowledge-intensive benchmark tasks while constraining the attention computation for each query to only use the top-k key-value pairs with the highest attention score (see Section 3.2 for algorithmic details). In these experiments, we use the same number of keys for each layer of the network and average the model performance over all tasks on the OpenLLM leaderboard (see Section 4.1 for task description). In Figure 2 we see that all three models saturate in performance by the 15th key, and the most recent (and most overtrained) variants of the model require only the 10 top keys to achieve the same performance as full-scale attentions. While Figure 3 indicates that there is some spreading of the attention mass for layer 1 of the network, this tail mass seems to be unnecessary for good performance on benchmarks.
Taken as a whole, these simple but motivating experiments evidence the basic sparsity assumptions on which our method is based. Empirically, we find that the attention scores are a native, reliable indicator as to which key-value pairs are most critical during any given inference step, and that very few keys are actually needed to perform accurate inference.
In the next section we introduce the technical details of our approach for using a vector database to retrieve the top-k most influential tokens, thereby disconnecting attention’s selection operations from the rest of the feed-forward computation. This critical separation enables the GPU to perform smaller matrix operations that fit into its limited memory whilst the full cached keys and values live in CPU memory and are searched using CPU compute. In the main experiments that follow, we showcase how our method alleviates both the computational and memory barriers that normally bottleneck long-context inference and unlocks million token contexts for small GPUs.
@@ 3 Methodology
In the following sections we elaborate on our method for reducing memory costs during inference time using top-k attention. Along the way, we more concretely define some minimal notation and terminology required to describe inference operations in transformer language models.
@@ 3.2 Top-k Attention: Accelerating Decoding
The $V_{\ell}[I]$ and vals are then moved to the GPU. Using these, we perform the final attention computation, combining both the weighted values from the context as well as the value vectors from previously generated tokens. This method gives us a peak GPU memory cost of $O(k)$ GPU memory cost, as opposed to $O(N)$. We find that $k$ can be chosen to be a small fraction (around 1% for most tasks) of $N$ while still recovering near-equivalent performance.
@@ 3.3 Prefilling a KV Cache at the Million Token Scale
For our largest experiments, we utilize Flash Attention and an H100 GPU to prefill the caches $N$=1M tokens of context. This allows us to generate exact prefilled KV caches with only small modifications to the network to accommodate the extreme memory requirements. The exact changes we made in order to prefill caches can be found in the Appendix and we employ a chunking strategy similar to (Gupta et al., 2021). We note that our experiments mimic an important use case of our method, namely a user with a large amount of documents that has access to only a limited amount of compute. In this case, the user can rent the required compute on the cloud for prefill once, then make as many queries as they want quickly with our method on their own hardware.
@@ 4 Evaluating Top-k Attention at Scale
We evaluate top-k to highlight the relationship between different values of k and performance. We find that models of various sizes and generations perform well even at small k. In general, we observe that a k equal to 2% of the total length of the context is sufficient to achieve 95% of the performance achieved with full, standard attention.
4.1 Benchmark Details
We analyze the effectiveness of top-k attention across three benchmarks: RULER, Open LLM Leaderboard v1, and AlpacaEval. Each of these benchmarks highlight a different aspect of top-k attention impact on LLM performance. RULER specifically tests long context abilities and we use Open LLM Leaderboard v1 and AlpacaEval together to examine how different model sizes and families perform under top-k attention. Measuring performance across these benchmarks presents a comprehensive understanding of how the top-k attention mechanism affects a model.
@@ 4.1 Benchmark Details
Open LLM Leaderboard Tasks
We investigate the performance of top-k on Open LLM Leaderboard tasks. In particular, we evaluate different models on various values of k on the average of MMLU (Hendrycks et al., 2020), ARC tasks both easy and challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), winogrande (Keisuke et al., 2019), OpenbookQA (Mihaylov et al., 2018), BoolQ (Clark et al., 2019), and PiQA (Bisk et al., 2020). For each task, we record the normalized accuracy when available; otherwise, we record accuracy. We report the average over tasks. We evaluate the following models on these benchmarks: Llama-1 (7B), Llama-2 (7B), Llama-3 (8B), Llama-3.1 (8B), Vicuna-v1.3 (7B), Llama-2 chat (7B), Llama-3 Instruct (8B), Llama-3.1 instruct (8B), Llama-3.2 1B instruct, and Llama-3.2 3B instruct (Touvron et al., 2023a; Zheng et al., 2023; Touvron et al., 2023b; Dubey et al., 2024; AI, 2024). The experiments are conducted using lm-eval-harness in a zero-shot setting (Gao et al., 2024).
@@
We benchmark top-k attention on AlpacaEval (Dubois et al., 2024). AlpacaEval $2.0$ requires a model to generate responses to $805$ queries. These responses are then compared by an LLM-as-a-Judge with GPT-4 Turbo responses generated from the same query set. The winrate percentage is the reported metric, and the LLM-as-a-Judge is GPT-4 Turbo.
4.2 Evaluation Results
Top-k Performance on RULER
We evaluate our method on RULER using GradientAI’s Llama-3-8B model that was trained using a context length of 262k tokens. For this model, we find that very small values of k are sufficient to recover near-baseline performance. At every context length evaluated, $95\%$ of the baseline performance can always be achieved with a k value of $1\%$ or less of the total length. In Table 1, at $k=2$, greater than $60\%$ performance is achieved at all context lengths. The performance on RULER improves as k increases. Nevertheless, even at a context length of $131$k tokens to achieve $\sim 98\%$ performance of the full attention only $12.5\%$ of the attention scores are required.
@@ 4.2 Evaluation Results
Interestingly, NIAH tasks have high success rates with little variation across choice of k. Question-answering tasks (from SQuAD (Rajpurkar et al., 2016) and HotpotQA (Yang et al., 2018)) have lower success rates with the 8B-size model, but also show low variation in success rates across different k. In our experiments, word-counting tasks (CWE and FWE) were the tasks most affected by the choice of k. This suggests that in practice, extra compute (larger k) should be allocated to these types of tasks to maintain performance. Table 2 shows the differences between each of these tasks in terms of what k-value is required to reach near-parity performance.
Top-k Performance on OpenLLM Leaderboard Tasks
We evaluate various models to find that top-k behavior exists regardless of instruction tuning, model size, or the number of tokens the model was trained on. Figure 6 left shows that regardless of the number of tokens on which the model was trained, all exhibit a similar trend with performance saturating at k values $<10$. When comparing Figure 6 left and center, we see that instruction models and pretrained base models exhibit similar behavior, saturating very quickly. Finally, in Figure 6 right, while different models achieve different performance maximums on the benchmark suite, we see that the effect of top-k is independent of model size.
@@
Figure 7 shows that the 95% performance with 2% of attention result holds on generation-intensive tasks as measured by AlpacaEval 2.0. The trend of small k values achieving near-baseline performance holds true across model sizes, and additional results showing performance across model generations can be found in the appendix (Figure 13).
Figure 7: k equivalent to 2% of context length is sufficient to achieve 95% of dense-attention performance on AlpacaEval 2.0, regardless of model size.
@@
Figure 8: One million token NIAH performance comparing cache eviction (Xiao et al., 2023) and top-k attention. The red cells show that attention sinks are incapable of retrieving tokens outside of the local window or early sink tokens. Top-k attention achieves 100% success with just $k=10$.
@@
To demonstrate the extreme scaling that top-k attention permits, we use our method to generate tokens conditioned on a context window of one million tokens. We choose RULER’s Needle In A Haystack task to use for the context, and we use GradientAI’s Llama-3-8B model that was trained out to a context length of 1M (Gradient Team, 2024). We run this experiment on a single GPU with a Faiss vector database prefilled with the contents of a KV cache. [^1] We use a variety of k values for this experiment find that k =1 is sufficient to solve Needle In A Haystack with a context length of 1 million tokens. We also compare our method against (Xiao et al., 2023) for needle in a haystack in Figure 8. Note that top-k attention has a 100% success rate regardless of where in the context the needle is hidden, which cannot be said for any available cache-eviction method.
4.4 Exploring Non-Uniform Choices of k
In this final section we report the results of our studies on the minimal k required for strong performance across different task types and on the impact of applying our top-k operator non-uniformly over the layers of the transformer.
Optimal k Across Task Types
Table 2 shows how the value of k necessary to achieve 95% of base model performance varies across the tasks in RULER. For the needle in a haystack task, only a tiny fraction of the entire context is necessary to achieve 95% of the base performance of the model, whereas nearly 9% is necessary in the case of Word Counting. Most of the tasks fall under the 1% line.
Table 2: Percent of attention scores required to reach 95% of dense attention performance for different categories of tasks from the RULER benchmark. k requirements were measured by performance on contexts from 8,000-128,000 tokens. All samples from Llama-3-8B.